GNU Parted User Manual

GNU Parted, version 3.2, 15 June 2014

Andrew Clausen clausen@gnu.org
Richard M. Kreuter kreuter@anduril.rutgers.edu

Leslie Patrick Polzer polzer@gnu.org

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999-2011 Free Software Foundation, Inc. Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, with no Front-Cover Texts, and with no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Short Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Using Parted . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3 Related information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 A Copying This Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 B This manual’s history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

1 Introduction

1.1 Overview of GNU Parted

GNU Parted is a program for creating and manipulating partition tables.

This documentation is written with the assumption that the reader has some understanding of partitioning and file systems.

GNU Parted was designed to minimize the chance of data loss. For example, it was designed to avoid data loss during interruptions (like power failure) and performs many safety checks. However, there could be bugs in GNU Parted, so you should back up your important files before running Parted.

    The GNU Parted homepage is http://www.gnu.org/software/parted. The library

and frontend themselves can be downloaded from ftp://ftp.gnu.org/gnu/parted. You can also find a listing of mailing lists, notes for contributing and more useful information on the web site.

Please send bug reports to bug-parted@gnu.org. When sending bug reports, please include the version of GNU Parted. Please include the output from these commands (for

disk /dev/hda):
     # parted /dev/hda print unit s print unit chs print

Feel free to ask for help on this list — just check that your question isn’t answered here first. If you don’t understand the documentation, please tell us, so we can explain it better. General philosophy is: if you need to ask for help, then something needs to be fixed so you (and others) don’t need to ask for help.

Also, we’d love to hear your ideas :-)

1.2 Software Required for the use of Parted

If you’re installing or compiling Parted yourself, you’ll need to have some other programs installed. If you are compiling Parted, you will need both the normal and devel packages of these programs installed:

• libuuid, part of the e2fsprogs package. If you don’t have this, you can get it from:

    http://web.mit.edu/tytso/www/linux/e2fsprogs.html

If you want to compile Parted and e2fsprogs, note that you will need to make install

    and make install-libs e2fsprogs.

• GNU Readline (optional), available from

    ftp://ftp.gnu.org/gnu/readline

If you are compiling Parted, and you don’t have readline, you can disable Parted’s readline support with the --disable-readline option for configure.

• GNU gettext (or compatible software) for compilation, if internationalisation support is desired.

    ftp://ftp.gnu.org/gnu/gettext

• libreiserfs, if you want reiserfs support:

    http://reiserfs.osdn.org.ua

Note that parted will automatically detect libreiserfs at runtime, and enable reiserfs support. libreiserfs is new, and hasn’t been widely tested yet.

1.3 Platforms on which GNU Parted runs

Hopefully, this list will grow a lot. If you do not have one of these platforms, then you can use a rescue disk and a static binary of GNU Parted.

GNU/Linux Linux versions 2.0 and up, on Alpha, x86 PCs, PC98, Macintosh PowerPC, Sun hardware.

GNU/Hurd

1.4 Terms of distribution for GNU Parted

GNU Parted is free software, covered by the GNU General Public License Version 2. This should have been included with the Parted distribution, in the COPYING file. If not, write to the Free Software Foundation, Inc., 51 Franklin St, Fifth Floor, Boston, MA 02110-1301, USA.

Libparted is considered part of GNU Parted. It is covered by the GNU General Public License. It is NOT released under the GNU Lesser General Public License (LGPL).

1.5 Building GNU Parted

If you want to compile GNU Parted, this is generally done with:

     $ ./configure
     $ make

However, there are a few options for configure:

--without-readline

turns off use of readline. This is useful for making rescue disks, etc., where few libraries are available.

--disable-debug

don’t include assertions

--disable-dynamic-loading

disables dynamic loading of some libraries (only libreiserfs for now, although we hope to expand this). Dynamic loading is useful because it allows you to reuse libparted shared libraries even when you don’t know if some libraries will be available. It has a small overhead (mainly linking with libdl), so it may be useful to disable it on bootdisks if you don’t need the flexibility.

--disable-nls

turns off native language support. This is useful for use with old versions of glibc, or a trimmed down version of glibc suitable for rescue disks.

--disable-shared

turns off shared libraries. This may be necessary for use with old versions of GNU libc, if you get a compile error about a “spilled register”. Also useful for boot/rescue disks.

--enable-discover-only

support only reading/probing (reduces size considerably)

--enable-mtrace

enable malloc() debugging

--enable-read-only

disable writing (for debugging)

1.5.1 Introduction

If you want to run GNU Parted on a machine without GNU/Linux installed, or you want to modify a root or boot partition, use GParted Live: http://gparted.sourceforge.net/

livecd.php.

2 Using Parted

2.1 Introduction to Partitioning

Unfortunately, partitioning your disk is rather complicated. This is because there are interactions between many different systems that need to be taken into consideration.

This manual used to introduce the reader to these systems and their working. This content has moved to the GNU Storage Guide.

2.2 Using GNU Parted

Parted has two modes: command line and interactive. Parted should always be started with:

     # parted device

where device is the hard disk device to edit. (If you’re lazy and omit the DEVICE argument, Parted will attempt to guess which device you want.)

In command line mode, this is followed by one or more commands. For example:

     # parted /dev/sda mklabel gpt mkpart P1 ext3 1MiB 8MiB

Options (like --help) can only be specified on the command line.

In interactive mode, commands are entered one at a time at a prompt, and modify the disk immediately. For example:

     (parted) mklabel gpt
     (parted) mkpart P1 ext3 1MiB 8MiB

Unambiguous abbreviations are allowed. For example, you can type “p” instead of “print”, and “u” instead of “units”. Commands can be typed either in English, or your native language (if your language has been translated). This may create ambiguities. Commands are case-insensitive.

Numbers indicating partition locations can be whole numbers or decimals. The suffix selects the unit, which may be one of those described in Section 2.4.14 [unit], page 11, except CHS and compact. If no suffix is given, then the default unit is assumed. Negative numbers count back from the end of the disk, with “-1s” indicating the sector at the end of the disk. Parted will compute sensible ranges for the locations you specify (e.g. a range of +/- 500 MB when you specify the location in “G”). Use the sector unit “s” to specify exact locations. With parted-2.4 and newer, IEC binary units like “MiB”, “GiB”, “TiB”, etc., specify exact locations as well. See [IEC binary units], page 12.

If you don’t give a parameter to a command, Parted will prompt you for it. For example:

     (parted) mklabel
     New disk label type? gpt

Parted will always warn you before doing something that is potentially dangerous, unless the command is one of those that is inherently dangerous (viz., rm, mklabel and mkpart). Since many partitioning systems have complicated constraints, Parted will usually do something slightly different to what you asked. (For example, create a partition starting at 10.352Mb, not 10.4Mb) If the calculated values differ too much, Parted will ask you for confirmation.

2.3 Command Line Options

When invoked from the command line, Parted supports the following syntax:

     # parted [option] device [command [argument]]

Available options and commands follow. For detailed explanations of the use of Parted commands, see Section 2.4 [Command explanations], page 5. Options begin with a hyphen, commands do not:

Options:

‘-h’

‘--help’ display a help message

‘-s’
‘--script’

never prompt the user

‘-a alignment-type’
‘--align alignment-type’

Set alignment for newly created partitions, valid alignment types are: none, cylinder, minimal and optimal.

‘-v’
‘--version’

display the version

2.4 Parted Session Commands

GNU Parted provides the following commands:

Note that after version 2.4, the following commands were removed: check, cp, mkfs, mkpartfs, move, resize.

2.4.1 align-check

[Command] align-check align-type n

Determine whether the starting sector of partition n meets the disk’s selected alignment criteria. align-type must be ‘minimal’, ‘optimal’ or an abbreviation. When in script mode, if the partition does not meet the alignment requirement, exit with status 1; otherwise (including on older kernels for which alignment data is not available), continue processing any remaining commands. Without --script, print either

     ‘N aligned’ or ‘N not aligned’.

Example:

          (parted) align-check minimal 1
          1 aligned

2.4.2 disk set

[Command] disk_set flag state

Changes a flag on the disk. A flag can be either “on” or “off”. Some or all of these flags will be available, depending on what disk label you are using:

‘pmbr_boot’

(GPT) - this flag enables the boot flag on the GPT’s protective MBR partition.

The disk’s flags are displayed by the print command on the "Disk Flags:" line. They are also output as the last field of the disk information in machine mode.

     (parted) disk_set pmbr_boot on

Set the PMBR’s boot flag.

2.4.3 help

[Command] help [command]

Prints general help, or help on command. Example:

     (parted) help mklabel

Print help for the mklabel command.

2.4.4 mklabel

[Command] mklabel label-type

Creates a new disk label, of type label-type. The new disk label will have no partitions. This command (normally) won’t technically destroy your data, but it will make it basically unusable, and you will need to use the rescue command (see Chapter 3 [Related information], page 14) to recover any partitions. Parted works on all partition tables.^{1} label-type must be one of these supported disk labels:

• bsd

• loop (raw disk access)

• gpt

• mac

• msdos

• pc98

• sun Example:

     (parted) mklabel msdos

Create an MS-DOS disk label. This is still the most common disk label for PCs.

2.4.5 mkpart

[Command] mkpart [part-type fs-type name] start end

Creates a new partition, without creating a new file system on that partition. This is useful for creating partitions for file systems (or LVM, etc.) that Parted doesn’t support. You may specify a file system type, to set the appropriate partition code in the partition table for the new partition. fs-type is required for data partitions (i.e., non-extended partitions). start and end are the offset from the beginning of the disk, that is, the “distance” from the start of the disk. part-type is one of ‘primary’, ‘extended’ or ‘logical’, and may be specified only with ‘msdos’ or ‘dvh’ partition tables. A name must be specified for a ‘gpt’ partition table. Neither part-type nor name may be used with a ‘sun’ partition table. fs-type must be one of these supported file systems:

1 Everyone seems to have a different word for “disk label” — these are all the same thing: partition

table, partition map.

• ext2

• fat16, fat32

• hfs, hfs+, hfsx

• linux-swap

• NTFS

• reiserfs

• ufs

• btrfs

For example, the following creates a logical partition that will contain an ext2 file system. The partition will start at the beginning of the disk, and end 692.1 megabytes into the disk.

     (parted) mkpart logical 0.0 692.1

Now, we will show how to partition a low-end flash device (“low-end”, as of 2011/2012). For such devices, you should use 4MiB-aligned partitions^{2}. This command creates a tiny place-holder partition at the beginning, and then uses all remaining space to create the partition you’ll actually use:

     $ parted -s /dev/sdX -- mklabel msdos \
         mkpart primary fat32 64s 4MiB \
         mkpart primary fat32 4MiB -1s

Note the use of ‘--’, to prevent the following ‘-1s’ last-sector indicator from being interpreted as an invalid command-line option. The above creates two empty partitions. The first is unaligned and tiny, with length less than 4MiB. The second partition starts precisely at the 4MiB mark and extends to the end of the device. The next step is typically to create a file system in the second partition:

     $ mkfs.vfat /dev/sdX2

2.4.6 name

[Command] name number name

Sets the name for the partition number (GPT, Mac, MIPS and PC98 only). The name can be placed in quotes. Example:

2 Cheap flash drives will be with us for a long time to come, and, for them, 1MiB alignment is not enough.

Use at least 4MiB-aligned partitions. For details, see Arnd Bergman’s article, http://http://lwn.

net/Articles/428584/ and its many comments.

(parted) name 2 'Secret Documents'

Set the name of partition 2 to ‘Secret Documents’.

2.4.7 print

[Command] print [number]

Displays the partition table on the device parted is editing, or detailed information about a particular partition. Example:

(parted) print
Disk geometry for /dev/hda: 0.000-2445.679 megabytes
Disk label type: msdos
Minor    Start       End     Type      Filesystem  Flags
1          0.031    945.000  primary   fat32       boot, lba
2        945.000   2358.562  primary   ext2
3       2358.562   2445.187  primary   linux-swap
(parted) print 1
Minor: 1
Flags: boot, lba
File System: fat32
Size:            945.000Mb (0%)
Minimum size:     84.361Mb (0%)
Maximum size:   2445.679Mb (100%)

2.4.8 quit

[Command] quit

Quits Parted. It is only after Parted exits that the Linux kernel knows about the changes Parted has made to the disks. However, the changes caused by typing your commands will probably be made to the disk immediately after typing a command. However, the operating system’s cache and the disk’s hardware cache may delay this.

2.4.9 rescue

[Command] rescue start end

Rescue a lost partition that used to be located approximately between start and end. If such a partition is found, Parted will ask you if you want to create a partition for it. This is useful if you accidently deleted a partition with parted’s rm command, for example. Example:

(parted) print
Disk geometry for /dev/hdc: 0.000-8063.507 megabytes
Disk label type: msdos
Minor    Start       End     Type      Filesystem  Flags
1          0.031   8056.032  primary   ext3
(parted) rm

Partition number? 1
(parted) print
Disk geometry for /dev/hdc: 0.000-8063.507 megabytes
Disk label type: msdos
Minor    Start       End     Type      Filesystem  Flags

OUCH! We deleted our ext3 partition!!! Parted comes to the rescue...

(parted) rescue
Start? 0
End? 8056
Information: A ext3 primary partition was found at 0.031MB ->
8056.030MB.  Do you want to add it to the partition table?
Yes/No/Cancel? y
(parted) print
Disk geometry for /dev/hdc: 0.000-8063.507 megabytes
Disk label type: msdos
Minor    Start       End     Type      Filesystem  Flags
1          0.031   8056.032  primary   ext3

It’s back! :)

2.4.10 resizepart

[Command] resizepart number end

Moves the end position of partition number. Note that this does not modify any filesystem present in the partition. If you wish to do this, you will need to use external tools, such as resize2fs.

When growing a partition you will want to grow the filesystem afterwards, but when shrinking, you need to shrink the filesystem before the partition.

2.4.11 rm

[Command] rm number

Removes the partition with number number. If you accidently delete a partition with this command, use mkpart to recover it. Also, you can use the gpart program (see Chapter 3 [Related information], page 14) to recover damaged disk labels.

Note for msdos disk labels: if you delete a logical partition, all logical partitions with a larger partition number will be renumbered. For example, if you delete a logical partition with a partition number of 6, then logical partitions that were number 7, 8 and 9 would be renumbered to 6, 7 and 8 respectively. This means, for example, that you have to update /etc/fstab on GNU/Linux systems.

Example:

(parted) rm 3

Remove partition 3.

2.4.12 select

[Command] select device

Selects the device, device, for Parted to edit. The device can be a Linux hard disk device, a partition, a software RAID device or LVM logical volume. Example:

(parted) select /dev/hdb

Select /dev/hdb (the slave device on the first ide controller on Linux) as the device to edit.

2.4.13 set

[Command] set number flag state

Changes a flag on the partition with number number. A flag can be either “on” or “off”. Some or all of these flags will be available, depending on what disk label you are using:

‘bios_grub’

(GPT) - Enable this to record that the selected partition is a GRUB BIOS partition.

‘legacy_boot’

(GPT) - this flag is used to tell special purpose software that the GPT partition may be bootable.

‘boot’ (Mac, MS-DOS, PC98) - should be enabled if you want to boot off the partition. The semantics vary between disk labels. For MS-DOS disk labels, only one partition can be bootable. If you are installing LILO on a partition that partition must be bootable. For PC98 disk labels, all ext2 partitions must be bootable (this is enforced by Parted).

‘msftdata’

(GPT) - This flag identifies partitions that contain Microsoft filesystems (NTFS or FAT). It may optionally be set on Linux filesystems to mimic the type of configuration created by parted 3.0 and earlier, in which a separate Linux filesystem type code was not available on GPT disks. This flag can only be removed within parted by replacing it with a competing flag, such as boot or msftres.

‘msftres’ (GPT) - This flag identifies a "Microsoft Reserved" partition, which is used by Windows on GPT disks. Note that this flag should not normally be set on Windows filesystem partitions (those that contain NTFS or FAT filesystems).

‘irst’ (MS-DOS, GPT) - this flag identifies an Intel Rapid Start Technology partition.

‘esp’ (MS-DOS, GPT) - this flag identifies a UEFI System Partition. On GPT it is an alias for boot.

‘lba’ (MS-DOS) - this flag can be enabled to tell MS DOS, MS Windows 9x and MS Windows ME based operating systems to use Linear (LBA) mode.

‘root’ (Mac) - this flag should be enabled if the partition is the root device to

be used by Linux.

‘swap’ (Mac) - this flag should be enabled if the partition is the swap device to

be used by Linux.

‘hidden’ (MS-DOS, PC98) - this flag can be enabled to hide partitions from Mi-

crosoft operating systems.

‘raid’ (MS-DOS) - this flag can be enabled to tell linux the partition is a software

RAID partition.

‘LVM’ (MS-DOS) - this flag can be enabled to tell linux the partition is a physical

volume.

‘PALO’ (MS-DOS) - this flag can be enabled so that the partition can be used by

the Linux/PA-RISC boot loader, palo.

‘PREP’ (MS-DOS, GPT) - this flag can be enabled so that the partition can be used as a PReP boot partition on PowerPC PReP or IBM RS6K/CHRP hardware.

‘DIAG’ (MS-DOS) - Enable this to indicate that a partition can be used as a

diagnostics / recovery partition.

The print command displays all enabled flags for each partition. Example:

     (parted) set 1 boot on

Set the ‘boot’ flag on partition 1.

2.4.14 unit

[Command] unit unit

Selects the current default unit that Parted will use to display locations and capacities on the disk and to interpret those given by the user if they are not suffixed by an unit. unit may be one of: ‘s’ sector (n bytes depending on the sector size, often 512) ‘B’ byte ‘KiB’ kibibyte (1024 bytes) ‘MiB’ mebibyte (1048576 bytes) ‘GiB’ gibibyte (1073741824 bytes) ‘TiB’ tebibyte (1099511627776 bytes) ‘kB’ kilobyte (1000 bytes) ‘MB’ megabyte (1000000 bytes) ‘GB’ gigabyte (1000000000 bytes) ‘TB’ terabyte (1000000000000 bytes)

‘%’ percentage of the device (between 0 and 100)

‘cyl’ cylinders (related to the BIOS CHS geometry)

‘chs’ cylinders, heads, sectors addressing (related to the BIOS CHS geometry)

‘compact’ This is a special unit that defaults to megabytes for input, and picks a unit that gives a compact human readable representation for output.

The default unit apply only for the output and when no unit is specified after an input number. Input numbers can be followed by an unit (without any space or other character between them), in which case this unit apply instead of the default unit for this particular number, but CHS and cylinder units are not supported as a suffix. If no suffix is given, then the default unit is assumed. Parted will compute sensible ranges for the locations you specify (e.g., a range of +/- 500 MB when you specify the location in “G”, and a range of +/- 500 KB when you specify the location in “M”) and will select the nearest location in this range from the one you wrote that satisfies constraints from both the operation, the filesystem being worked on, the disk label, other partitions and so on. Use the sector unit “s” to specify exact locations (if they do not satisfy all constraints, Parted will ask you for the nearest solution). Note that negative numbers count back from the end of the disk, with “-1s” pointing to the last sector of the disk.

Note that as of parted-2.4, when you specify start and/or end values using IEC binary units like “MiB”, “GiB”, “TiB”, etc., parted treats those values as exact, and equivalent to the same number specified in bytes (i.e., with the “B” suffix), in that it provides no “helpful” range of sloppiness. Contrast that with a partition start request of “4GB”, which may actually resolve to some sector up to 500MB before or after that point. Thus, when creating a partition, you should prefer to specify units of bytes (“B”), sectors (“s”), or IEC binary units like “MiB”, but not “MB”, “GB”, etc.

Example:

     (parted) unit compact
     (parted) print
     Disk geometry for /dev/hda: 0kB - 123GB
     Disk label type: msdos
     Number  Start   End     Size    Type      File system  Flags
     1       32kB    1078MB  1077MB  primary   reiserfs     boot
     2       1078MB  2155MB  1078MB  primary   linux-swap
     3       2155MB  123GB   121GB   extended
     5       2155MB  7452MB  5297MB  logical   reiserfs

(parted) unit chs print
Disk geometry for /dev/hda: 0,0,0 - 14946,225,62
BIOS cylinder,head,sector geometry: 14946,255,63.  Each cylinder
is 8225kB.
Disk label type: msdos
Number  Start       End         Type      File system  Flags
1       0,1,0       130,254,62  primary   reiserfs     boot
2       131,0,0     261,254,62  primary   linux-swap
3       262,0,0     14945,254,62 extended
5       262,2,0     905,254,62  logical   reiserfs
(parted) unit mb print
Disk geometry for /dev/hda: 0MB - 122942MB
Disk label type: msdos
Number  Start   End     Size    Type      File system  Flags
1       0MB     1078MB  1077MB  primary   reiserfs     boot
2       1078MB  2155MB  1078MB  primary   linux-swap
3       2155MB  122935MB 120780MB extended
5       2155MB  7452MB  5297MB  logical   reiserfs

3 Related information

If you want to find out more information, please see the GNU Parted web site.

These files in the Parted distribution contain further information:

• ABOUT-NLS - information about using Native Language Support, and the Free Translation Project.

• AUTHORS - who wrote what.

• ChangeLog - record of changes made to Parted.

• COPYING - the GNU General Public License, the terms under which GNU Parted may be distributed.

• COPYING.DOC - the GNU Free Documentation Licence, the term under which Parted’s documentation may be distributed.

• INSTALL — how to compile and install Parted, and most other free software

Appendix A Copying This Manual

A.1 GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text. The “publisher” means any person or entity that distributes copies of the Document to the public. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING

You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3.

You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY

If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects.

If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages.

If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public.

It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS

You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version:

A. Use in the Title Page (and on the covers, if any) a title distinct from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Appendix B This manual’s history

This manual was based on the file USER included in GNU Parted version 1.4.22 source distribution. The GNU Parted source distribution is available at ftp.gnu.org/gnu/

parted.

Initial Texinfo formatting by Richard M. Kreuter, 2002.

Maintainance by Andrew Clausen from 2002 to 2005 and by Leslie P. Polzer from July 2005 onwards.

This manual is distributed under the GNU Free Documentation License, version 1.1 or later, at your discretion, any later version published by the Free Software Foundation; with no Invariant Sections, with no Front-Cover Texts, and with no Back-Cover Texts. See Appendix A [Copying This Manual], page 15, for details.

Index 24 Index 25

Index

A

align-check, command description . . . . . . . . . . . . . . . 5

B

bugs, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 building parted . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

C

command description, align-check . . . . . . . . . . . . . . . 5 command description, disk set . . . . . . . . . . . . . . . . . . 5 command description, help. . . . . . . . . . . . . . . . . . . . . . 6 command description, mkindex . . . . . . . . . . . . . . . . . 6 command description, mkpart. . . . . . . . . . . . . . . . . . . 6 command description, name. . . . . . . . . . . . . . . . . . . . . 7 command description, print . . . . . . . . . . . . . . . . . . . . . 8 command description, quit . . . . . . . . . . . . . . . . . . . . . . 8 command description, rescue . . . . . . . . . . . . . . . . . . . . 8 command description, resizepart . . . . . . . . . . . . . . . . 9 command description, rm . . . . . . . . . . . . . . . . . . . . . . . 9 command description, select . . . . . . . . . . . . . . . . . . . 10 command description, set . . . . . . . . . . . . . . . . . . . . . . 10 command description, unit. . . . . . . . . . . . . . . . . . . . . 11 command syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 commands, detailed listing . . . . . . . . . . . . . . . . . . . . . . 5 commands, overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 compiling parted. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 contacting developers . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

D

description of parted . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 detailed command listing . . . . . . . . . . . . . . . . . . . . . . . 5 disk set, command description . . . . . . . . . . . . . . . . . . 5

E

e2fsprogs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

F

FDL, GNU Free Documentation License . . . . . . . 15 further reading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

G

gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 gnu gpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 gpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

H

help, command description. . . . . . . . . . . . . . . . . . . . . . 6 history of this manual . . . . . . . . . . . . . . . . . . . . . . . . . 23

I

invocation options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

L

libuuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 license terms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

M

mklabel, command description . . . . . . . . . . . . . . . . . . 6 mkpart, command description. . . . . . . . . . . . . . . . . . . 6 modes of use. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

N

name, command description. . . . . . . . . . . . . . . . . . . . . 7

O

options at invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 overview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

P

parted description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 partitioning overview . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 platforms, supported. . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 print, command description . . . . . . . . . . . . . . . . . . . . . 8

Q

quit, command description . . . . . . . . . . . . . . . . . . . . . . 8

R

readline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 related documentation . . . . . . . . . . . . . . . . . . . . . . . . . 14 reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 required software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 rescue, command description . . . . . . . . . . . . . . . . . . . . 8 resizepart, command description . . . . . . . . . . . . . . . . 9 rm, command description . . . . . . . . . . . . . . . . . . . . . . . 9

S

select, command description . . . . . . . . . . . . . . . . . . . 10 set, command description . . . . . . . . . . . . . . . . . . . . . . 10 software dependencies. . . . . . . . . . . . . . . . . . . . . . . . . . . 1 supported platforms . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

T

terms of distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

U

unit, command description. . . . . . . . . . . . . . . . . . . . . 11